the scale on a map of Virginia shows that 1 cm represents 30 miles. the actual distance from richmond, VA to Washington, DC is 110 miles. one the map, how many cm are between the 2 cities. round your answer to the nearest tenth.
Answer by stanbon(75887)

the scale on a map of Virginia shows that 1 cm represents 30 miles. the actual distance from richmond, VA to Washington, DC is 110 miles. one the map, how many cm are between the 2 cities. round your answer to the nearest tenth.
---
Use a proportion:
xcm/110 mi = 1cm/30 miles
---
x = 110/30 = 11/3 = 3 2/3 cm
==================
Cheers,
Stan H.